### 08 Mar 2026 v1

# On a conjecture concerning the property of chromatic polynomials with negative variable

Abstract
Let $G$ be a graph of order $n$ and $P(G,x)$ be the chromatic polynomial of $G$. Dong, Ge, Gong, Ning, Ouyang, and Tay (J. Graph Theory 96 (2021) 343) conjectured that $\frac{d^{k}}{dx^{k}}\bigl(\ln[(-1)^{n}P(G,x)]\bigr)<0$ holds for all $k\geq 2$ and $x\in(-\infty,0)$. We prove this conjecture for all $k\geq 2$ and $x\leq-10\Delta k$, in which $\Delta$ is the maximum degree of $G$.
Keywords: chromatic polynomial; higher derivative; root.
Mathematics Subject Classification (2020): 05C31.
1 Introduction
All graphs considered in this paper are simple and connected. For any graph $G=(V,E)$ and any positive integer $x$, a proper $x$-coloring of $G$ is a mapping $c:V\rightarrow\{1,\ldots,x\}$ such that $c(u)\neq c(v)$ whenever $uv\in E$. The chromatic number $\chi(G)$ is the least $x$ such that $G$ has a proper $x$-coloring. In 1912, Birkhoff [2] introduced the chromatic polynomial $P(G,x)$ of $G$ to count the number of distinct proper $x$-colorings of $G$, i.e.,
$P(G,x)=|\{c:V\rightarrow\{1,\ldots,x\}\mid c\text{ is proper}\}|.$
It was hoped that the chromatic polynomial could be used to solve the four color conjecture of that time, because there is a clear relationship between chromatic polynomials and chromatic numbers, i.e., $\chi(G)$ is the smallest positive integer such that $P(G,\chi(G))\neq 0$. Although this approach for the four color problem has not been successful so far, chromatic polynomials attracted widespread attention and have become an active topic nowadays. For a book devoted to the topic, see Dong, Koh and Teo [4].
By the principle of inclusion and exclusion, we have
$P(G,x)=\sum_{E^{\prime}\subset E}(-1)^{|E^{\prime}|}x^{\kappa(E^{\prime})},$
where $\kappa(E^{\prime})$ denotes the number of connected components in $E^{\prime}$. It is clear that $P(G,x)$ is a polynomial in $x$. In 1932, Whitney [11] interpreted the coefficients of $P(G,x)$ by introducing the notion of broken cycles. Let $\eta:E\to\{1,2,\dots,|E|\}$ be a bijection, for any cycle $C$ in $G$, the path CeC-e is called a broken cycle of $G$ if $e$ is the edge on $C$ with $\eta(e)\leq\eta(e^{\prime})$ for every edge $e^{\prime}$ on $C$.
Theorem 1.1 ([11]).
Let $G$ be a graph of order $n$ and $\eta:E\to\{1,2,\dots,|E|\}$ be a bijection. Then,
$P(G,x)=\sum_{i=1}^{n}(-1)^{n-i}a_{i}(G)x^{i},$
where $a_{i}(G)$ is the number of spanning subgraphs of $G$ with nin-i edges which do not contain broken cycles.
Let $x$ be an arbitrary complex number in $P(G,x)$, and $\Delta(G)$ (or simply $\Delta$) be the maximum degree of a graph $G$. The distribution of the roots of $P(G,x)=0$ and evaluations of $P(G,x)$ at some specific points are two of the many research problems related to chromatic polynomials. We now introduce the results concerning the bound on the roots of $P(G,x)=0$ that are needed in this note.
Theorem 1.2 ([10]).
There exists a constant $K$ such that for every graph $G$, $P(G,x)\neq 0$ for all $x\in\mathbb{C}$ with $|x|\geq K\Delta$.
Sokal [10] showed the constant $K\approx 7.97$, Fernández and Procacci [6] improved to $K\approx 6.91$, Jenssen, Patel, Regts [8] recently showed $K\leq 5.94$.
Denote $P^{(i)}(G,x)$ the $i$-th derivative of $P(G,x)$. Recently, Bernardi and Nadeau [1] gave a nice combinatorial interpretation of $P^{(i)}(G,-j)$ for any nonnegative integers $i$ and $j$ in terms of acyclic orientations. When $i=0$ and $j=0$, their result recovers the classical interpretations due to Stanley [9] and to Greene and Zaslavsky [7] respectively. In this note, we study the property of $(\ln[(-1)^{n}P(G,x)])^{(k)}$ for $k\geq 1$ and $x<0$.
In 2020, Dong, Ge, Gong, Ning, Ouyang, and Tay [3] studied a parameter called the mean size of a broken-cycle-free spanning subgraph of a graph $G$ of order $n$, that is $\epsilon(G)=\sum\limits_{i=1}^{n}(n-i)a_{i}(G)/\sum\limits_{i=1}^{n}a_{i}(G)$, where $a_{i}(G)$ is the number as defined in Theorem $1.$ They proved that $\epsilon(T_{n})<\epsilon(G)<\epsilon(K_{n})$ holds for any graph $G$ of order $n$ which is neither the complete graph $K_{n}$ nor a tree $T_{n}$ of order $n$. In their proof, they introduced a function $\epsilon(G,x)=P^{\prime}(G,x)/P(G,x)$. One can deduce that $\epsilon(G)=n+\epsilon(G,-1)$ holds for any graph $G$ of order $n$. Thus, for any graphs $G$ and $H$ of the same order, $\epsilon(G)<\epsilon(H)$ is equivalent to $\epsilon(G,-1)<\epsilon(H,-1)$. In [3], the authors proved that for any graph $G$ of order $n$ which is neither the complete graph nor a tree, $\epsilon(T_{n},x)<\epsilon(G,x)<\epsilon(K_{n},x)$ holds for all $x<0$. Then $\epsilon(T_{n})<\epsilon(G)<\epsilon(K_{n})$ follows.
It is clear that for any graph $G$ of order $n$,
$\epsilon(G,x)=\frac{P^{\prime}(G,x)}{P(G,x)}=\frac{d}{dx}(\ln[(-1)^{n}P(G,x)])<0$
holds for all $x<0$. Because for $x<0$, $(-1)^{n}P(G,x)>0$ and $(-1)^{n}P^{\prime}(G,x)<0$. In [3], the authors conjectured that this property holds for higher derivatives of the function $\ln[(-1)^{n}P(G,x)]$ for $x<0$.
Conjecture 1.3 ([3]).
Let $G$ be a graph of order $n$. Then
$\frac{d^{k}}{dx^{k}}(\ln[(-1)^{n}P(G,x)])<0$
holds for all $k\geq 2$ and $x\in(-\infty,0)$.
We focus on Conjecture 1.3 in this note, and prove the conjecture for all $k\geq 2$ and $x\leq-10\Delta k$. For $k=2$ or 33, we have the following results.
Lemma 1.4.
Let $G$ be a graph of order $n$ and $K$ be the constant in Theorem 1.2. Then
$\frac{d^{2}}{dx^{2}}(\ln[(-1)^{n}P(G,x)])<0$
holds for $x<-2K\Delta$, and
$\frac{d^{3}}{dx^{3}}(\ln[(-1)^{n}P(G,x)])<0$
holds for $x<-(\sqrt{3}+1)K\Delta$.
Proof.
Suppose that the real roots of $P(G,x)=0$ are $x_{1},\ldots,x_{i}$ and the complex roots are $a_{1}\pm ib_{1},\ldots,a_{j}\pm ib_{j}$, then $i+2j=n$, and
$P(G,x)=(x-x_{1})\cdots(x-x_{i})(x^{2}-2a_{1}x+a_{1}^{2}+b_{1}^{2})\cdots(x^{2}-2a_{j}x+a_{j}^{2}+b_{j}^{2}).$
By computing, we have
$\frac{d^{2}}{dx^{2}}(\ln[(-1)^{n}P(G,x)])=\sum_{t=1}^{i}\frac{-1}{(x-x_{t})^{2}}+\sum_{t=1}^{j}\frac{2b_{t}^{2}-2(x-a_{t})^{2}}{\big((x-a_{t})^{2}+b_{t}^{2}\big)^{2}}$
and
$\frac{d^{3}}{dx^{3}}(\ln[(-1)^{n}P(G,x)])=\sum_{t=1}^{i}\frac{2}{(x-x_{t})^{3}}+\sum_{t=1}^{j}\frac{4(x-a_{t})\big((x-a_{t})^{2}-3b_{t}^{2}\big)}{\big((x-a_{t})^{2}+b_{t}^{2}\big)^{3}}.$
It is well known that the $P(G,x)=0$ has no negative real roots. Hence, for $x<0$, $\frac{-1}{(x-x_{t})^{2}}<0$ and $\frac{2}{(x-x_{t})^{3}}<0$. When $x<-(|a_{t}|+|b_{t}|)$, $\frac{2b_{t}^{2}-2(x-a_{t})^{2}}{\big((x-a_{t})^{2}+b_{t}^{2}\big)^{2}}<0$; when $x<-(|a_{t}|+|\sqrt{3}b_{t}|)$, $\frac{4(x-a_{t})\big((x-a_{t})^{2}-3b_{t}^{2}\big)}{\big((x-a_{t})^{2}+b_{t}^{2}\big)^{3}}<0$.
From Theorem 1.2, $|a_{t}+ib_{t}|\leq K\Delta$. Then $x<-2K\Delta$ implies $x<-(|a_{t}|+|b_{t}|)$, $x<-(\sqrt{3}+1)K\Delta$ implies $x<-(|a_{t}|+|\sqrt{3}b_{t}|)$. The lemma follows. ∎
Combining Lemma 1.4 and $K\leq 5.94$, we get that $(\ln[(-1)^{n}P(G,x)])^{\prime\prime}<0$ for $x<-12\Delta$, and $(\ln[(-1)^{n}P(G,x)])^{\prime\prime\prime}<0$ for $x<-6(\sqrt{3}+1)\Delta$. When $k$ becomes larger, the method in the above lemma is no longer efficient. In Section 2, we present a different approach to solve this problem.
2 Main results
For a graph $G$ of order $n$, let $\alpha_{1},\ldots,\alpha_{n}$ be the roots of $P(G,x)=0$ and
$c_{i}=-\dfrac{1}{i}\sum\limits_{j=1}^{n}\alpha_{j}^{i}.$
Fadnavis [5] deduced the exact values of $c_{1}$ and $c_{2}$.
Lemma 2.1 ([5]).
Let $|T|$ be the number of triangles in $G$, then $c_{1}=-|E|$ and $c_{2}=-|T|-|E|/2.$
Let $K$ be the constant in Theorem 1.2. Using Taylor expansion, one can deduce the following result.
Theorem 2.2 ([5]).
For a graph $G$ of order $n$ and $|x|>K\Delta$, we have
$\ln\frac{P(G,x)}{x^{n}}=\sum_{i=1}^{\infty}c_{i}x^{-i}.$
Lemma 2.3.
For $x\leq q$, $q<-K\Delta$, we have
(i) $\sum\limits_{i=1}^{\infty}c_{i}x^{-i}$ converges uniformly.
(ii) $\sum\limits_{i=1}^{\infty}(c_{i}x^{-i})^{(k)},~~k=1,2,\ldots$, converges uniformly.
(iii) Both $\sum\limits_{\begin{subarray}{c}i=1\\ \text{$i$ is odd}\end{subarray}}^{\infty}\big((\frac{K\Delta}{x})^{i}\big)^{(k)}$ and $\sum\limits_{\begin{subarray}{c}i=1\\ \text{$i$ is even}\end{subarray}}^{\infty}\big((\frac{K\Delta}{x})^{i}\big)^{(k)}$,  $k=1,2,\ldots$, converge uniformly.
Proof.
The proofs of the three results are similar, so we only present the proof of conclusion (ii). From the definition of $c_{i}$ and Theorem 1.2, we have $|c_{i}|\leq\frac{1}{i}\sum\limits_{j=1}^{n}|\alpha_{j}|^{i}\leq\frac{n}{i}(K\Delta)^{i}$. Then, for $x\leq q$, $q<-K\Delta$, we have
$|(c_{i}x^{-i})^{(k)}|=|c_{i}\dfrac{(-1)^{k}i(i+1)\cdots(i+k-1)}{x^{i+k}}|\leq\dfrac{n}{i}(K\Delta)^{i}\dfrac{i(i+1)\cdots(i+k-1)}{(-q)^{i+k}}.$
And the positive term series $\sum\limits_{i=1}^{\infty}\dfrac{(i+1)\cdots(i+k-1)}{(-q)^{k}}(\dfrac{K\Delta}{-q})^{i}$ converges. From Weierstrass M-test, $\sum\limits_{i=1}^{\infty}(c_{i}x^{-i})^{(k)}$, $x\leq q$, $q<-K\Delta$ converges uniformly. ∎
From Lemma 2.3 and properties of uniform convergence series, we have
[…]
for $k\geq 0$ ($k=0$ refers to the series itself), and the term-by-term differentiation is valid.
Because $K\leq 5.94$, $|c_{i}|\leq\frac{n}{i}(K\Delta)^{i}<\frac{n}{i}(6\Delta)^{i}$. And
$(c_{i}x^{-i})^{(k)}=c_{i}\frac{(-1)^{k}i(i+1)\cdots(i+k-1)}{x^{i+k}}.$
For $x\leq q$, $q<-6\Delta$ and $k\geq 2$, when $i$ is odd, $\frac{(-1)^{k}}{x^{i+k}}<0$,
$(c_{i}x^{-i})^{(k)}\leq\dfrac{-n(6\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=n(6\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)};$
when $i$ is even, $\frac{(-1)^{k}}{x^{i+k}}>0$,
$(c_{i}x^{-i})^{(k)}\leq\dfrac{n(6\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=-n(6\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)}.$
Hence
[…]
and
[…]
Lemma 2.4.
For a graph $G$ of order $n$, $k\geq 1$ and $x<0$,
$(\ln[(-1)^{n}P(G,x)])^{(k)}=(\ln\frac{P(G,x)}{x^{n}})^{(k)}+\frac{(k-1)!(-1)^{k-1}n}{x^{k}}.$
Proof.
If $x<0$, then $\ln\frac{P(G,x)}{x^{n}}=\ln(-1)^{n}P(G,x)-\ln(-1)^{n}x^{n}$. Because $(\ln(-1)^{n}x^{n})^{(k)}=\frac{(k-1)!(-1)^{k-1}n}{x^{k}}$, the lemma follows. ∎
Theorem 2.5.
For a graph $G$ of order $n$ and $k\geq 2$,
$(\ln[(-1)^{n}P(G,x)])^{(k)}<0,$
holds for all $x<-10\Delta k$.
Proof.
If $n=1$, then $P(G,x)=x$, $(\ln(-x))^{(k)}=\frac{(-1)^{k-1}(k-1)!}{x^{k}}<0$ holds for all $x<0$. We now assume $n\geq 2$.
From Theorem 2.2 and Lemma 2.4,
$(\ln[(-1)^{n}P(G,x)])^{(k)}=(\sum\limits_{i=1}^{\infty}c_{i}x^{-i})^{(k)}+\frac{(k-1)!(-1)^{k-1}n}{x^{k}}.$ (4)
Now we consider $(\sum\limits_{i=1}^{\infty}c_{i}x^{-i})^{(k)}$. From (1),
[…]
From Lemma 2.1, $c_{1}=-|E|$ and $c_{2}=-|T|-|E|/2.$ By computing,
$(\frac{c_{1}}{x})^{(k)}+(\frac{c_{2}}{x^{2}})^{(k)}$ $=$ $-|E|\frac{(-1)^{k}\cdot k!}{x^{k+1}}-(|T|+\frac{|E|}{2})\frac{(-1)^{k}\cdot(k+1)!}{x^{k+2}}$
Because $G$ is a connected graph with order $n\geq 2$, and the maximum degree is $\Delta$, we have $\frac{n}{2}\leq|E|\leq\frac{\Delta n}{2}$. Then
$(\frac{c_{1}}{x})^{(k)}+(\frac{c_{2}}{x^{2}})^{(k)}$ $\leq$ $-\frac{\Delta n}{2}\frac{(-1)^{k}k!}{x^{k+1}}-\frac{n}{4}\frac{(-1)^{k}(k+1)!}{x^{k+2}}.$ (5)
By (2)(3), some computations and simplifications, we obtain
$(\sum\limits_{\begin{subarray}{c}i=3\\ \text{$i$ is odd}\end{subarray}}^{\infty}c_{i}x^{-i})^{(k)}+(\sum\limits_{\begin{subarray}{c}i=3\\ \text{$i$ is even}\end{subarray}}^{\infty}c_{i}x^{-i})^{(k)}$ (6)
$\leq$ $\frac{n}{6\Delta}(\sum_{\begin{subarray}{c}i=3\\ i\text{ is odd}\end{subarray}}^{\infty}(\frac{6\Delta}{x})^{i+1})^{(k-1)}-\frac{n}{6\Delta}(\sum_{\begin{subarray}{c}i=4\\ i\text{ is even}\end{subarray}}^{\infty}(\frac{6\Delta}{x})^{i+1})^{(k-1)}$
$=$ $\frac{n}{6\Delta}(\frac{(6\Delta)^{4}}{x^{3}(x+6\Delta)})^{(k-1)}=n(\frac{1}{x}-\frac{6\Delta}{x^{2}}+\frac{(6\Delta)^{2}}{x^{3}}-\frac{1}{x+6\Delta})^{(k-1)}$
$=$ $n(-1)^{k-1}\Big(\frac{(k-1)!}{x^{k}}-6\Delta\frac{k!}{x^{k+1}}+(6\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+6\Delta)^{k}}\Big)$
Combining (4), (5) and (6), we have
$(\ln[(-1)^{n}P(G,x)])^{(k)}$
$\leq$ $-\frac{\Delta n}{2}\frac{(-1)^{k}k!}{x^{k+1}}-\frac{n}{4}\frac{(-1)^{k}(k+1)!}{x^{k+2}}$
$+n(-1)^{k-1}\Big(\frac{2(k-1)!}{x^{k}}-6\Delta\frac{k!}{x^{k+1}}+(6\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+6\Delta)^{k}}\Big)$
$=$ $n(-1)^{k-1}(k-1)!\frac{(x+6\Delta)^{k}\big(8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)\big)-4x^{k+2}}{4x^{k+2}(x+6\Delta)^{k}}$
To prove $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ for $x<-10\Delta k$, it is enough to verify the following inequality for $x<-10\Delta k$,
$(-1)^{k-1}(x+6\Delta)^{k}\big(8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)\big)<(-1)^{k-1}4x^{k+2}.$ (7)
We denote $f_{1}(x)=8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)$. Because $f_{1}(x)>0$ and $(-1)^{k-1}x^{k}<0$, inequality (7) is equivalent to
$\frac{(-1)^{k-1}(x+6\Delta)^{k}}{(-1)^{k-1}x^{k}}>\frac{4x^{2}}{f_{1}(x)},$
that is
$(1+\frac{6\Delta}{x})^{k}>\frac{4x^{2}}{f_{1}(x)}.$
By using Bernoulli’s Inequality, we have $(1+\frac{6\Delta}{x})^{k}\geq 1+\frac{6\Delta k}{x}.$ We let
$1+\frac{6\Delta k}{x}>\frac{4x^{2}}{f_{1}(x)}$
which is equivalent to
$xf_{1}(x)+6k\Delta f_{1}(x)-4x^{3}<0.$
Let $f(x)=xf_{1}(x)+6k\Delta f_{1}(x)-4x^{3}.$ We can check that when $x\leq-6\Delta k$, $f^{\prime}(x)>0$, $f(x)$ is increasing strictly. By computing, we have $f(-10\Delta k)<0$. Hence when $x\leq-10\Delta k$, $f(x)<0$, $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ follows. The proof is completed. ∎
Acknowledgements
The author is supported by National Natural Science Foundation of China (No. 12371350).

### 16 Mar 2026 v2

@@
-On a conjecture concerning the property of chromatic polynomials with negative variable
+On a conjecture concerning the property of chromatic polynomials with negative variables
 Abstract
@@
 Graph Theory 96 (2021) 343) conjectured that $\frac{d^{k}}{dx^{k}}\bigl(\ln[(-1)^{n}P(G,x)]\bigr)<0$ holds for all $k\geq 2$ and $x\in(-\infty,0)$.
-We prove this conjecture for all $k\geq 2$ and $x\leq-10\Delta k$, in which $\Delta$ is the maximum degree of $G$.
+We prove this conjecture for all $k\geq 2$ and $x\leq-6.66\Delta k$, in which $\Delta$ is the maximum degree of $G$.
 Keywords: chromatic polynomial; higher derivative; root.
@@
 Theorem 1.2 ([11]).
-There exists a constant $K$ such that for every graph $G$, $P(G,x)\neq 0$ for all $x\in\mathbb{C}$ with $|x|\geq K\Delta$.
-Sokal [10] showed the constant $K\approx 7.97$, Fernández and Procacci [6] improved to $K\approx 6.91$, Jenssen, Patel, Regts [8] recently showed $K\leq 5.94$.
+There exists a constant $K$ such that for every graph $G$, $P(G,x)\neq 0$ for all $x\in\mathbb{C}$ with $|x|>K\Delta$.
+Sokal [11] showed the constant $K\leq 7.97$, Fernández and Procacci [7] improved to $K\leq 6.91$, Jenssen, Patel, Regts [9] further improved to $K\leq 5.94$.
+Very recently, Bencs and Regts [3] established a better bound $K\leq 4.25$.
 Denote $P^{(i)}(G,x)$ the $i$-th derivative of $P(G,x)$.
@@
 for $k\geq 0$ ($k=0$ refers to the series itself), and the term-by-term differentiation is valid.
-Because $K\leq 5.94$, $|c_{i}|\leq\frac{n}{i}(K\Delta)^{i}<\frac{n}{i}(6\Delta)^{i}$. And
+Because $K\leq 4.25$, $|c_{i}|\leq\frac{n}{i}(K\Delta)^{i}\leq\frac{n}{i}(4.25\Delta)^{i}$. And
 $(c_{i}x^{-i})^{(k)}=c_{i}\frac{(-1)^{k}i(i+1)\cdots(i+k-1)}{x^{i+k}}.$
-For $x\leq q$, $q<-6\Delta$ and $k\geq 2$, when $i$ is odd, $\frac{(-1)^{k}}{x^{i+k}}<0$,
-$(c_{i}x^{-i})^{(k)}\leq\dfrac{-n(6\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=n(6\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)};$
+For $x\leq q$, $q<-4.25\Delta$ and $k\geq 2$, when $i$ is odd, $\frac{(-1)^{k}}{x^{i+k}}<0$,
+$(c_{i}x^{-i})^{(k)}\leq\dfrac{-n(4.25\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=n(4.25\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)};$
 when $i$ is even, $\frac{(-1)^{k}}{x^{i+k}}>0$,
-$(c_{i}x^{-i})^{(k)}\leq\dfrac{n(6\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=-n(6\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)}.$
+$(c_{i}x^{-i})^{(k)}\leq\dfrac{n(4.25\Delta)^{i}(-1)^{k}(i+1)\cdots(i+k-1)}{x^{i+k}}=-n(4.25\Delta)^{i}(\frac{1}{x^{i+1}})^{(k-1)}.$
 Hence
@@
 $(\ln[(-1)^{n}P(G,x)])^{(k)}<0,$
-holds for all $x<-10\Delta k$.
+holds for all $x\leq-6.66\Delta k$.
 Proof.
@@
 $(\sum\limits_{\begin{subarray}{c}i=3\\ \text{$i$ is odd}\end{subarray}}^{\infty}c_{i}x^{-i})^{(k)}+(\sum\limits_{\begin{subarray}{c}i=3\\ \text{$i$ is even}\end{subarray}}^{\infty}c_{i}x^{-i})^{(k)}$ (6)
-$\leq$ $\frac{n}{6\Delta}(\sum_{\begin{subarray}{c}i=3\\ i\text{ is odd}\end{subarray}}^{\infty}(\frac{6\Delta}{x})^{i+1})^{(k-1)}-\frac{n}{6\Delta}(\sum_{\begin{subarray}{c}i=4\\ i\text{ is even}\end{subarray}}^{\infty}(\frac{6\Delta}{x})^{i+1})^{(k-1)}$
-$=$ $\frac{n}{6\Delta}(\frac{(6\Delta)^{4}}{x^{3}(x+6\Delta)})^{(k-1)}=n(\frac{1}{x}-\frac{6\Delta}{x^{2}}+\frac{(6\Delta)^{2}}{x^{3}}-\frac{1}{x+6\Delta})^{(k-1)}$
-$=$ $n(-1)^{k-1}\Big(\frac{(k-1)!}{x^{k}}-6\Delta\frac{k!}{x^{k+1}}+(6\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+6\Delta)^{k}}\Big)$
+$\leq$ $\frac{n}{4.25\Delta}(\sum_{\begin{subarray}{c}i=3\\ i\text{ is odd}\end{subarray}}^{\infty}(\frac{4.25\Delta}{x})^{i+1})^{(k-1)}-\frac{n}{4.25\Delta}(\sum_{\begin{subarray}{c}i=4\\ i\text{ is even}\end{subarray}}^{\infty}(\frac{4.25\Delta}{x})^{i+1})^{(k-1)}$
+$=$ $\frac{n}{4.25\Delta}(\frac{(4.25\Delta)^{4}}{x^{3}(x+4.25\Delta)})^{(k-1)}=n(\frac{1}{x}-\frac{4.25\Delta}{x^{2}}+\frac{(4.25\Delta)^{2}}{x^{3}}-\frac{1}{x+4.25\Delta})^{(k-1)}$
+$=$ $n(-1)^{k-1}\Big(\frac{(k-1)!}{x^{k}}-4.25\Delta\frac{k!}{x^{k+1}}+(4.25\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+4.25\Delta)^{k}}\Big)$
 Combining (4), (5) and (6), we have
@@
 $\leq$ $-\frac{\Delta n}{2}\frac{(-1)^{k}k!}{x^{k+1}}-\frac{n}{4}\frac{(-1)^{k}(k+1)!}{x^{k+2}}$
-$+n(-1)^{k-1}\Big(\frac{2(k-1)!}{x^{k}}-6\Delta\frac{k!}{x^{k+1}}+(6\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+6\Delta)^{k}}\Big)$
-$=$ $n(-1)^{k-1}(k-1)!\frac{(x+6\Delta)^{k}\big(8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)\big)-4x^{k+2}}{4x^{k+2}(x+6\Delta)^{k}}$
-To prove $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ for $x<-10\Delta k$, it is enough to verify the following inequality for $x<-10\Delta k$,
-$(-1)^{k-1}(x+6\Delta)^{k}\big(8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)\big)<(-1)^{k-1}4x^{k+2}.$ (7)
-We denote $f_{1}(x)=8x^{2}-22x\Delta k+k(k+1)+2(6\Delta)^{2}k(k+1)$. Because $f_{1}(x)>0$ and $(-1)^{k-1}x^{k}<0$, inequality (7) is equivalent to
-$\frac{(-1)^{k-1}(x+6\Delta)^{k}}{(-1)^{k-1}x^{k}}>\frac{4x^{2}}{f_{1}(x)},$
+$+n(-1)^{k-1}\Big(\frac{2(k-1)!}{x^{k}}-4.25\Delta\frac{k!}{x^{k+1}}+(4.25\Delta)^{2}\frac{(k+1)!}{2x^{k+2}}-\frac{(k-1)!}{(x+4.25\Delta)^{k}}\Big)$
+$=$ $n(-1)^{k-1}(k-1)!\frac{(x+4.25\Delta)^{k}\big(8x^{2}-15x\Delta k+k(k+1)+2(4.25\Delta)^{2}k(k+1)\big)-4x^{k+2}}{4x^{k+2}(x+4.25\Delta)^{k}}$
+To prove $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ for $x\leq-6.66\Delta k$, it is enough to verify the following inequality for $x\leq-6.66\Delta k$,
+$(-1)^{k-1}(x+4.25\Delta)^{k}\big(8x^{2}-15x\Delta k+k(k+1)+2(4.25\Delta)^{2}k(k+1)\big)<(-1)^{k-1}4x^{k+2}.$ (7)
+We denote $f_{1}(x)=8x^{2}-15x\Delta k+k(k+1)+2(4.25\Delta)^{2}k(k+1)$. Because $f_{1}(x)>0$ and $(-1)^{k-1}x^{k}<0$, inequality (7) is equivalent to
+$\frac{(-1)^{k-1}(x+4.25\Delta)^{k}}{(-1)^{k-1}x^{k}}>\frac{4x^{2}}{f_{1}(x)},$
 that is
-$(1+\frac{6\Delta}{x})^{k}>\frac{4x^{2}}{f_{1}(x)}.$
-By using Bernoulli’s Inequality, we have $(1+\frac{6\Delta}{x})^{k}\geq 1+\frac{6\Delta k}{x}.$ We let
-$1+\frac{6\Delta k}{x}>\frac{4x^{2}}{f_{1}(x)}$
+$(1+\frac{4.25\Delta}{x})^{k}>\frac{4x^{2}}{f_{1}(x)}.$
+By using Bernoulli’s Inequality, we have $(1+\frac{4.25\Delta}{x})^{k}\geq 1+\frac{4.25\Delta k}{x},$ when $x<-4.25\Delta k$. We let
+$1+\frac{4.25\Delta k}{x}>\frac{4x^{2}}{f_{1}(x)}$
 which is equivalent to
-$xf_{1}(x)+6k\Delta f_{1}(x)-4x^{3}<0.$
-Let $f(x)=xf_{1}(x)+6k\Delta f_{1}(x)-4x^{3}.$ We can check that when $x\leq-6\Delta k$, $f^{\prime}(x)>0$, $f(x)$ is increasing strictly.
-By computing, we have $f(-10\Delta k)<0$.
-Hence when $x\leq-10\Delta k$, $f(x)<0$, $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ follows.
-The proof is completed.
+$xf_{1}(x)+4.25k\Delta f_{1}(x)-4x^{3}<0.$
+Let $f(x)=xf_{1}(x)+4.25k\Delta f_{1}(x)-4x^{3}.$ We can check that when $x<-4.25\Delta k$, $f^{\prime}(x)>0$, $f(x)$ is increasing strictly.
+By computing, we have $f(-6.66\Delta k)<0$.
+Hence when $x\leq-6.66\Delta k$, $f(x)<0$, $(\ln[(-1)^{n}P(G,x)])^{(k)}<0$ follows.
+The proof is complete.
 ∎
+We note that to obtain inequalities (2) and (3) required for the proof of Theorem 2.5, we need to estimate the value of $c_{i}$.
+In this note, we estimate $c_{i}$ by using $|c_{i}|\leq\frac{n}{i}(K\Delta)^{i}$.
+Thus, a new bound on $K$ gives a new bound on $c_{i}$, and a smaller $K$ yields a better bound.
 Acknowledgements
-The author is supported by National Natural Science Foundation of China (No. 12371350).
+The author thanks Guus Regts for informing her of reference [3], which provides the best bound on $K$ to date and is very helpful for the present note.
+The author is supported by National Natural Science Foundation of China (No.
+12371350).

### 19 Mar 2026 v3

@@
 holds for all $k\geq 2$ and $x\in(-\infty,0)$.
-We focus on Conjecture 1.3 in this note, and prove the conjecture for all $k\geq 2$ and $x\leq-10\Delta k$. For $k=2$ or 33, we have the following results.
+We focus on Conjecture 1.3 in this note, and prove the conjecture for all $k\geq 2$ and $x\leq-6.66\Delta k$. For $k=2$ or 33, we have the following results.
 Lemma 1.4.
@@
 From Theorem 1.2, $|a_{t}+ib_{t}|\leq K\Delta$. Then $x<-2K\Delta$ implies $x<-(|a_{t}|+|b_{t}|)$, $x<-(\sqrt{3}+1)K\Delta$ implies $x<-(|a_{t}|+|\sqrt{3}b_{t}|)$. The lemma follows. ∎
-Combining Lemma 1.4 and $K\leq 5.94$, we get that $(\ln[(-1)^{n}P(G,x)])^{\prime\prime}<0$ for $x<-12\Delta$, and $(\ln[(-1)^{n}P(G,x)])^{\prime\prime\prime}<0$ for $x<-6(\sqrt{3}+1)\Delta$.
+Combining Lemma 1.4 and $K\leq 4.25$, we get that $(\ln[(-1)^{n}P(G,x)])^{\prime\prime}<0$ for $x<-8.5\Delta$, and $(\ln[(-1)^{n}P(G,x)])^{\prime\prime\prime}<0$ for $x<-4.25(\sqrt{3}+1)\Delta$.
 When $k$ becomes larger, the method in the above lemma is no longer efficient.